Sodium hydroxide, NaOH, is a strong base that is sometimes used as a cleaning agent since it can be used to dissolve proteins, grease, oils, and fats.
Rank the following in increasing order of their molarity.
Rank from lowest to highest. To rank items as equivalent, overlap them.
a. 63.0 g of NaOH dissolved in 1.0 L of water
b. 6.625 g of NaOH dissolved in 250. mL of water
c. 8.0 g of NaOH dissolved in 500. mL of water
d. .530 g of NaOH dissolved in 100 mL of water
1
Expert's answer
2015-11-05T10:56:27-0500
The molarity equals:
c = n/V, where n - number of moles, V - volume of water in L.
n = m/M, where M - is a molar mass, m - is a mass in g.

C(NaOH) = m(NaOH)/(M(NaOH) V(water))
C (a) = 63/40 1 = 1.575 mol/L
C (b) = 6.625/40 0.25 = 0.663 mol/L
C (c) = 8/40 0.5 = 0.4 mol/L
C (d) = 0.53/40 0.1 = 0.133 mol/L

Answer:
d - c - b - a